# Through the LLM Looking Glass: A Socratic Self-Assessment of Donkeys, Elephants, and Markets

### v1

@@ Abstract
While detecting and avoiding bias in LLM-generated text is becoming increasingly important, media bias often remains subtle and subjective, making it particularly difficult to identify and mitigate. In this study, we assess media bias in LLM-generated content and LLMs’ ability to detect subtle ideological bias. We conduct this evaluation using two datasets, PoliGen and EconoLex, covering political and economic discourse, respectively. We evaluate eight widely used LLMs by prompting them to generate articles and analyze their ideological preferences via self-assessment. By using self-assessment, the study aims to directly measure the models’ biases rather than relying on external interpretations, thereby minimizing subjective judgments about media bias. Our results reveal a consistent preference of Democratic over Republican positions across all models. Conversely, in economic topics, biases vary among Western LLMs, while those developed in China lean more strongly toward socialism.
Through the LLM Looking Glass: A Socratic Self-Assessment of Donkeys, Elephants, and Markets
@@ 1 Introduction
Figure 1: Left: Democratic vs. Republican preferences in self-generated articles on political topics. Right: Socialist vs. Capitalist preferences in self-generated articles on economical titles.
@@
Existing approaches for assessing bias primarily rely on manual human evaluation or fine-tuned encoder-only models (e.g., reward models). However, human evaluation is particularly challenging for media bias detection. Beyond being expensive, media bias is often subtle and subjective Spinde et al. (2022), and human annotators may themselves hold biases, making objective assessment difficult. Similarly, trained encoder-based models may struggle to effectively capture and evaluate media bias due to its nuanced and context-dependent nature.
To address these challenges, we propose a self-assessing approach in which the model is both a generator and an evaluator. Using a Socratic method, [^1] Which has shown promising results in existing research He et al. (2024). the model generates biased content and selects its preferred response. Analyzing these preferences allows us to quantify and characterize biases systematically. Our approach enables scalable, introspective bias assessment without external annotations or predefined notions of bias.
We present a systematic study of the degree of bias in eight widely used LLMs across various political and economic topics, followed by further analysis of LLMs’ integrity and agentic behavior. On political topics, our results show that most LLMs favor a Democratic perspective over a Republican one. In economic topics, Western-developed models remain relatively neutral, whereas models developed in China lean more strongly toward a socialist perspective, complementing the findings of Buyl et al. (2025). Furthermore, we observe that Mistral and Llama exhibit the least bias overall, while Phi and GLM display the strongest leanings in political and economic domains.
@@ 2 Related Work
Political Bias in LLMs.
Recent studies show that large language models (LLMs) exhibit various biases, notably political bias Bender et al. (2021); Weidinger et al. (2022); Bommasani et al. (2021). Some works demonstrate how slight prompt changes can shift ideological stances Bang et al. (2024), while instruction-tuned LLMs often reinforce existing user ideologies Hernandes and Corsi (2024). Beyond politics, studies on fairness, gender, and religious biases Vig et al. (2020); Felkner et al. (2023); Abid et al. (2021); Reif and Schwartz (2024) reveal emergent patterns linked to instruction tuning Itzhak et al. (2024).
Despite these advances, many methods focus on media bias without fully leveraging LLMs Spinde et al. (2023); Horych et al. (2025), or target broad fairness rather than specific political issues Motoki et al. (2024); Gehman et al. (2020). Such gaps underscore the need for deeper evaluations of political bias across diverse topics and expanded use of self-assessment mechanisms Bang et al. (2024); Lin et al. (2024).
LLM Self-Assessment in Media Bias. LLM-based media bias detection remains challenging due to alignment with dominant narratives in training data Spinde et al. (2023); Horych et al. (2025). Expert-labeled datasets (e.g., BABE) capture bias partially but lack broader ideological coverage Spinde et al. (2023). Various methods—such as fine-tuning on bias indicators Lin et al. (2024), diverse model ensembling Horych et al. (2025), and self-reflective prompting Bang et al. (2024); Schick et al. (2021)—aim to mitigate these limitations. Notably, instruction-tuned models often display more consistent ideological drift Trhlík and Stenetorp (2024). Esiobu et al. (2023) propose ROBBIE, a framework for robust bias evaluation, by applying continuous monitoring.
@@ 4 Methodology
In the first stage, we focused on generating articles across different ideological perspectives. Each model was prompted to generate five articles per topic/title, alternating between Democratic, Republican, and neutral perspectives for PoliGen (§3) and Socialist, Capitalist, and neutral perspectives for EconoLex (§3). To ensure a balanced and systematic generation process, we employed predefined system and user prompt combinations. The system and user prompt configurations are detailed in Table 2 for PoliGen and in Table 3 for EconoLex. The generated articles were then used as inputs for Stage 2, where models selected their preferred article on the same topic, generated from different ideological perspectives.
@@
In the second stage, we conducted an evaluation to assess model preferences as indicators of bias. Each LLM was presented with two or three articles on the same topic that it had generated in Stage 1 and instructed to choose out of these two or three articles the one it preferred. See Table 6 for the details on the prompts we used. To standardize preference indication, the prompts are desigend to elicit comparative judgments between the articles. At this stage (Stage 2), the LLM was unaware of each article’s perspective. Thus, a preference for a specific perspective over others can be interpreted as the model’s leaning toward that perspective.
@@ 5 Experiment Setup
We selected eight state-of-the-art LLMs from different families for a comprehensive analysis. Table 1 provides their details.
@@
In Stage 1 (§4), we used each model’s default generation parameters for both datasets. Each prompt followed the corresponding model’s chat template, incorporating a system message and user input. Articles were generated with a maximum of $512512$ tokens. For preference indication (§4), we constrained the output to a maximum of $1010$ tokens and employed greedy decoding to ensure deterministic responses. The output was parsed using regular expressions to extract the selected number. If a model refused to respond during the preference step or the output could not be processed, the sample was skipped. This occurred in fewer than one percent of samples for a few models.
@@
Table 4: Comparison of Models Across Political and Economic Dimensions. For each column, the largest value is in bold, and the second-largest value is underlined.
Two Way Political Three Way Political Two Way Economic Three Way Economic Model Democrat Republican Democrat Republican Neutral Socialist Capitalist Socialist Capitalist Neutral CMD R+ 0.52 0.48 0.35 0.33 0.32 0.45 0.55 0.26 0.34 0.40 R1 0.54 0.46 0.34 0.33 0.33 0.48 0.52 0.30 0.36 0.33 Qwen 0.59 0.41 0.45 0.21 0.34 0.67 0.33 0.35 0.27 0.38 Mistral 0.59 0.41 0.50 0.22 0.28 0.49 0.51 0.41 0.32 0.27 Phi-4 0.76 0.24 0.44 0.17 0.39 0.49 0.51 0.34 0.33 0.33 GLM 0.74 0.26 0.48 0.15 0.37 0.59 0.41 0.45 0.28 0.27 Gemma 0.67 0.33 0.49 0.12 0.39 0.54 0.46 0.31 0.29 0.40 Llama 0.54 0.46 0.49 0.17 0.34 0.47 0.52 0.27 0.39 0.33
@@ 6 Experiment results
Table 4 reports two-way and three-way preference outcomes for political and economic topics across the evaluated LLMs. Articles are generated using the first and second rows of Table 2 and Table 3, while preference indication is evaluated using and prompts 1 and 4 from Table 6.
In political topics, most models exhibit a Democratic leaning, with Phi, GLM, and Gemma showing the strongest tilt, while CMD R+, R1, and Llama remain relatively balanced. A large portion of Phi’s training data comes from the GPT series Abdin et al. (2024), potentially influencing its bias. In economic topics, most LLMs lean marginally to one side, but Qwen and GLM—both developed in China—strongly favor socialist perspectives, possibly reflecting regional ideologies. These results are complementary to previous work suggesting that language models often reflect the ideologies of their creators (Buyl et al., 2025).
While one might expect a heavy preference shift towards a neutral perspective upon its addition to the available options, in political classification the preferences barely shifts, while economic classification yields a more balanced socialist-capitalist-neutral split. This contrast likely stems from clearer political binaries (i.e., democratic vs. republican) versus broader economic viewpoints. Future work could explore whether these biases arise from lexical choices or content-related factors, as categorized in the media bias taxonomy Spinde et al. (2023).
@@
As we illustrate in Figure 2, using an LLM as a biased agent generally results in less bias than when a user requests biased content, except for the case of Mistral on political topics. Specifically, GLM and Phi (political topics) and GLM and Qwen (economic topics) act notably fairer in biased-agent mode. Given that system and user roles are introduced during post-training (supervised fine-tuning and RLHF), we speculate that richer and more varied content in user role data may have increased the LLM’s sensitivity to biased content requests in user prompts. Interestingly, CMD R+, one of the least biased models, exhibits minimal fluctuation across different roles, possibly due to more effective safety tuning. Eventually, we observed less fluctuation in economic topics than in political ones in this study.
Figure 2: Comparison of LLMs’ preferences when the article is generated by a biased agent versus a biased user. Left: Political topics. Right: Economic topics.
Figure 3: Comparison of LLMs’ behavior when asked to choose their preferred article versus when asked to identify which article is less biased. Left: Political topics. Right: Economic topics.

### v2

# Through the LLM Looking Glass: A Socratic Probing of Donkeys, Elephants, and Markets

@@ Abstract
While detecting and avoiding bias in LLM-generated text is becoming increasingly important, media bias often remains subtle and subjective, making it particularly difficult to identify and mitigate. In this study, we assess media bias in LLM-generated content and LLMs’ ability to detect subtle ideological bias. We conduct this evaluation using two datasets, PoliGen and EconoLex, covering political and economic discourse, respectively. We evaluate seven widely used LLMs by prompting them to generate articles and analyze their ideological preferences via Socratic probing. By using our self-contained Socratic approach, the study aims to directly measure the models’ biases rather than relying on external interpretations, thereby minimizing subjective judgments about media bias. Our results reveal a consistent preference of Democratic over Republican positions across all models. Conversely, in economic topics, biases vary among Western LLMs, while those developed in China lean more strongly toward socialism.
Through the LLM Looking Glass: A Socratic Probing of Donkeys, Elephants, and Markets
@@ 1 Introduction
Figure 1: We first ask each LLM to generate a document $d$ with a particular preference (e.g., “Democratic”) and then ask it to assess $d$’s preference without including any information about the generation process. Results of this generation and self-assessment method are shown for eight LLMs. Left: LLMs prefer Democratic over Republican positions in self-generated articles on political topics. Right: While there is no consistent trend for Western models, Chinese models prefer Socialist over Capitalist positions in self-generated articles on economical titles.
@@
Existing approaches for assessing bias primarily rely on manual human evaluation or fine-tuned encoder-only models Horych et al. (2025); Spinde et al. (2022). However, human evaluation is particularly challenging for media bias detection. Beyond being expensive, media bias is often subtle and subjective Spinde et al. (2021), and human annotators may themselves hold biases, making objective assessment difficult. Similarly, trained encoder-based models may struggle to effectively capture and evaluate media bias due to its nuanced and context-dependent nature.
To address these challenges, we propose a self-contained approach in which the model is both a generator and an evaluator. Using a Socratic method, [^1] Which has shown promising results in existing research He et al. (2024). the model is presented with a sequence of prompts designed to gradually elicit its underlying ideological leaning. First, the model is prompted to generate biased content and then to select its preferred response (from among the generated content). Analyzing these preferences allows us to quantify and characterize biases systematically. Our approach, while significantly more affordable than human annotation, enables scalable, introspective bias assessment without external annotations or predefined notions of bias.
We present a systematic study of the degree of bias in seven widely used LLMs across various political and economic topics, followed by further analysis of LLMs’ integrity and agentic behavior. Given the proprietary nature of many popular commercial models, whose training data, alignment methods, and safety interventions are undisclosed, we deliberately focus on open-source models in our experiments. This ensures transparency, reproducibility, and comparability in our evaluations. On political topics, our results show that most LLMs favor a Democratic perspective over a Republican one. In economic topics, Western-developed models remain relatively neutral, whereas models developed in China lean more strongly toward a socialist perspective, complementing the findings of Buyl et al. (2025). Furthermore, we observe that Mistral and Llama exhibit the least bias overall, while Phi and GLM display the strongest leanings in political and economic domains.
To validate our Socratic probing approach, we conduct a human annotation study to evaluate both the generation and evaluation components of the method. First, to assess the quality of the bias-inducing prompts and the resulting generated content, human annotators evaluate a subset of the generated articles to verify whether the intended ideological bias is indeed present. Annotators achieve a $75\%$ accuracy in identifying the correct bias type, indicating that the generation process reliably produces distinguishable biased content. Second, to assess the model’s understanding of bias, we prompt the models to identify the most neutral text among a set of biased alternatives. We report and analyze the extent to which model judgments align with human annotations, providing insight into their internal representations of bias and neutrality.
@@ 2 Related Work
Political and Ideological Bias in Language Models
Recent studies have shown that large language models (LLMs) exhibit various forms of bias, including political, gender, and religious biases Bender et al. (2021); Weidinger et al. (2022); Bommasani et al. (2021); Vig et al. (2020); Felkner et al. (2023); Abid et al. (2021); Reif and Schwartz (2024). A growing body of work has specifically focused on political bias, investigating how LLMs may align with particular ideological stances.
Several studies demonstrate that small changes in prompts can significantly shift a model’s political leaning Bang et al. (2024), while instruction-tuned LLMs tend to reinforce users’ pre-existing ideologies Hernandes and Corsi (2024). Lin et al. (2024) introduce IndiVec, a method for capturing subtle ideological shifts resulting from minor lexical variations, and Trhlík and Stenetorp (2024) find that instruction-tuned models more strongly align with ideological narratives than their base counterparts.
Beyond politics, other dimensions of social bias have also been studied. For example, Itzhak et al. (2024) find that instruction tuning can amplify pre-existing tendencies in LLMs, while Abid et al. (2021) highlight systematic anti-Muslim bias. Broader efforts such as (Dhamala et al., 2021) aim to quantify open-ended bias through dataset development and targeted metrics.
Despite these advances, existing approaches often suffer from several limitations: (1) a predominant focus on media bias detection using encoder-only models without fully utilizing the introspective capabilities of generative LLMs Spinde et al. (2023); Horych et al. (2025), (2) an emphasis on general fairness rather than specific ideological or political leanings Motoki et al. (2024); Gehman et al. (2020), (3) evaluations restricted to narrow topics, which may miss more nuanced ideological variance Bang et al. (2024); Hernandes and Corsi (2024); Trhlík and Stenetorp (2024), and (4) a lack of self-contained mechanisms that enable models to critically examine their own outputs Lin et al. (2024).
LLM-Based Self-Assessment and Media Bias Detection
Media bias detection with LLMs remains a challenging task due to their training on data that may align with dominant ideological narratives Spinde et al. (2023); Horych et al. (2025). Traditional bias detection approaches have relied on expert-labeled datasets such as BABE Spinde et al. (2023), which, while useful, often lack the ideological breadth and granularity needed to evaluate subtle bias manifestations.
To overcome these limitations, recent work has explored LLM-centric techniques for media bias evaluation. Lin et al. (2024) fine-tune models on ideological indicators to enhance sensitivity to biased framing, while Bang et al. (2024) use contextualized prompting to guide self-reflective assessments. Schick et al. (2021) propose a dynamic self-diagnosis and self-debiasing framework, allowing models to identify and mitigate their biases during inference.
Further, ensemble approaches have been used to dilute the bias of individual models while preserving the broader contextual understanding LLMs offer Horych et al. (2025). Trhlík and Stenetorp (2024) analyze ideological consistency across 56,700 synthetic articles, showing that instruction-tuned models exhibit more predictable ideological alignment. Esiobu et al. (2023) introduce ROBBIE, a robust evaluation framework that monitors ideological drift over time, providing a benchmark for long-term bias assessment.
Our work builds upon these insights by proposing a Socratic probing method in which LLMs act as both content generators and evaluators in a Socratic framework. This approach addresses the limitations of previous studies by enabling introspective, scalable bias quantification without reliance on external annotations or narrow topic coverage, as further detailed in Section 5.
3 Media Bias and Framing
Media bias refers to systematic patterns in how news outlets frame events, select facts, and use language to privilege certain perspectives over others Entman (2007); McQuail and Deuze (2020). Unlike demographic or representational biases embedded within language models, media bias is inherently tied to the socio-political context of news reporting and is observable only within the outputs of specific outlets. It is expressed not only through explicit opinions or partisan narratives but also through more subtle mechanisms such as the prioritization of certain themes, the use of emotionally charged language, or the omission of alternative perspectives.
A clear illustration of media bias can be seen in the coverage of the Respect for Marriage Act Vallejo et al. (2023). In this case, left-leaning outlets framed the legislation as a historic win for LGBTQ+ rights, emphasizing bipartisan support and presenting celebratory imagery. In contrast, right-leaning outlets framed the event around political contestation and judicial overreach, focusing on potential constitutional challenges. This example highlights how media bias often operates through framing choices rather than overt factual distortions, subtly guiding audience interpretations.
In our study, we build upon this understanding of framing bias by evaluating the capacity of large language models to generate ideologically framed outputs that mirror such nuanced forms of bias. Although the models were prompted to produce articles from a range of ideological perspectives—including Democratic, Republican, Socialist, Capitalist, and Neutral—we illustrate the phenomenon here using an example contrasting Socialist and Capitalist framings (see Figure 2). These examples are drawn from our experiments, which are detailed in Section 6 . Given the same headline, "How to Focus in a Manic Market," the Socialist-framed article critiques the manic market as a symptom of capitalist systems prioritizing speculative profits over societal welfare, adopting a cautionary tone and highlighting systemic risks to ordinary people. In contrast, the Capitalist-framed article embraces market volatility as a source of opportunity, framing it as an inherent feature of dynamic economies and encouraging investors to capitalize on the chaos. These examples demonstrate how LLM-generated outputs can reflect framing bias through shifts in tone, emphasis, and value prioritization, paralleling how media outlets present the same events through different ideological lenses (see Figure 2).
Prompt: Write an article on the following title from the perspective of a Socialist journalist. Title: How to Focus in a Manic Market LLM-generated article: […] In today’s global economy, financial markets often resemble a rollercoaster ride, characterized by extreme volatility and unpredictability. This manic market behavior can be disorienting, especially for those who are not part of the financial elite. From a perspective centered on economic justice, it is crucial to understand the underlying forces driving these market fluctuations and to focus on strategies that prioritize collective well-being over individual profit. […] The manic market is a product of a system that prioritizes profit maximization and speculative investment over stable, long-term growth. This encourages risky behavior, as financial institutions and investors chase short-term gains, often at the expense of economic stability. The result is a market prone to bubbles and crashes, leaving ordinary people vulnerable to economic shocks.
Prompt: Write an article on the following title from the perspective of a Capitalist journalist. Title: How to Focus in a Manic Market LLM-generated article: […] In the ever-evolving landscape of global finance, markets can often swing from calm to chaotic with startling speed. The recent surge in market volatility has left many investors and business leaders grappling with uncertainty. However, for those embracing market principles, these manic markets present not just challenges, but opportunities. Here’s how to maintain focus and capitalize on the chaos. First and foremost, understanding that markets are inherently dynamic is crucial. Volatility is not an anomaly but a natural part of the economic cycle. Businesses thrive on change, and a manic market is a testament to the fluidity of supply and demand. Embrace this dynamism as a catalyst for innovation and growth.
Figure 2: LLM-generated articles with contrasting ideological framings. Both articles share the same headline but differ in framing emphasis. The left article critiques market instability as systemic failure, focusing on collective welfare and economic risks. The right article frames volatility as an opportunity for growth, emphasizing entrepreneurial resilience. Key phrases illustrating framing differences are bolded for clarity. Text excerpts are abridged for illustration.
@@ 5 Methodology
In the first stage, we focused on generating articles across different ideological perspectives. Each model was prompted to generate articles per topic/title, alternating between Democratic, Republican, and neutral perspectives for PoliGen (§4) and Socialist, Capitalist, and neutral perspectives for EconoLex (§4). To ensure a balanced and systematic generation process, we employed predefined system and user prompt combinations. The system and user prompt configurations are detailed in Table 2 for PoliGen and in Table 3 for EconoLex.
@@
In the second stage, we conducted an evaluation to assess model preferences as indicators of bias. The generated articles from Stage 1 were used as inputs for this stage. Given two (excluding neutral) or three (including neutral) articles on the same topic generated from different ideological perspectives, models were asked to identify their preferred article. See Table 8 for the details on the prompts we used. To standardize preference indication, the prompts are designed to elicit comparative judgments between the articles. At this stage (Stage 2), the LLM was unaware of each article’s perspective, i.e., the LLM was not informed about the State 1 generation process that produced the article. Thus, a preference for a specific perspective over others can be interpreted as the model’s leaning toward that perspective.
@@ 6 Experiment Setup
We selected seven state-of-the-art LLMs from different families for a comprehensive analysis. Table 1 provides their details.
@@
In Stage 1 (§5), we used each model’s default generation parameters for both datasets. Each prompt followed the corresponding model’s chat template, incorporating a system message and user input. Articles were generated with a maximum of $512$ tokens. For preference indication (§5), we constrained the output to a maximum of $10$ tokens and employed greedy decoding to ensure deterministic responses. The output was parsed using regular expressions to extract the selected number. If a model refused to respond during the preference step or the output could not be processed, the sample was skipped. This occurred in fewer than one percent of samples for a few models.
@@
Table 4: Comparison of Models Across Political and Economic Dimensions. “Two Way Political” uses row 1 of Table 2. “Three Way Political” uses rows 1 and 2 of Table 2. “Two Way Economic” uses row 1 of Table 3. “Three Way Economic” uses rows 1 and 2 of Table 3. For each column, the largest value is in bold, and the second-largest value is underlined.
Two Way Political Three Way Political Two Way Economic Three Way Economic Model Democrat Republican Democrat Republican Neutral Socialist Capitalist Socialist Capitalist Neutral CMD R+ 0.52 0.48 0.35 0.33 0.32 0.45 0.55 0.26 0.34 0.40 Qwen 0.59 0.41 0.45 0.21 0.34 0.67 0.33 0.35 0.27 0.38 Mistral 0.59 0.41 0.50 0.22 0.28 0.49 0.51 0.41 0.32 0.27 Phi-4 0.76 0.24 0.44 0.17 0.39 0.49 0.51 0.34 0.33 0.33 GLM 0.74 0.26 0.48 0.15 0.37 0.59 0.41 0.45 0.28 0.27 Gemma 0.67 0.33 0.49 0.12 0.39 0.54 0.46 0.31 0.29 0.40 Llama 0.54 0.46 0.49 0.17 0.34 0.47 0.52 0.27 0.39 0.33
@@ 7 Experiment results
Table 4 reports two-way and three-way preference outcomes for political and economic topics across the evaluated LLMs. Articles are generated using the first and second rows of Table 2 and Table 3, while preference indication is evaluated using prompts 1 and 2 from Table 8.
In political topics, most models exhibit a Democratic leaning, with Phi, GLM, and Gemma showing the strongest tilt, while CMD R+, and Llama remain relatively balanced. A large portion of Phi’s training data comes from the GPT series Abdin et al. (2024), potentially influencing its bias. In economic topics, most LLMs lean marginally to one side, but Qwen and GLM—both developed in China—strongly favor socialist perspectives, possibly reflecting regional ideologies. These results are complementary to previous work suggesting that language models often reflect the ideologies of their creators (Buyl et al., 2025).
While one might expect a heavy preference shift towards a neutral perspective upon its addition to the available options, in political classification the preference barely shifts, while economic classification yields a more balanced socialist-capitalist-neutral split. This contrast likely stems from clearer political binaries (i.e., democratic vs. republican) versus broader economic viewpoints. Future work could explore whether these biases arise from lexical choices or content-related factors, as categorized in the media bias taxonomy Spinde et al. (2023).
@@
As we illustrate in Figure 3, using an LLM as a biased agent generally results in less bias than when a user requests biased content, except for the case of Mistral on political topics. Specifically, GLM and Phi (political topics) and GLM and Qwen (economic topics) act notably fairer in biased-agent mode. Given that system and user roles are introduced during post-training (supervised fine-tuning and RLHF), we speculate that richer and more varied content in user role data may have increased the LLM’s sensitivity to biased content requests in user prompts. Interestingly, CMD R+, one of the least biased models, exhibits minimal fluctuation across different roles, possibly due to more effective safety tuning. Finally, we observed less fluctuation in economic topics than in political ones in this study.
Figure 3: Comparison of LLMs’ preferences when the article is generated by a biased agent (rows 3 of Tables 2 and 3) versus a biased user (rows 1 of Tables 2 and 3). Left: Political topics. Right: Economic topics.
Figure 4: Comparison of LLMs’ behavior when asked to choose their preferred article (prompt 1 in Table 8) versus when asked to identify which article is less biased (prompt 3 in Table 8). Left: Political topics. Right: Economic topics.
@@
7.3 Human Annotation Study
Annotator Profile and Procedure
To evaluate the alignment between model outputs and their intended ideological prompts, we conducted an annotation study using model-generated articles based on titles from the previously introduced EconoLex dataset (Section 4). Its ideologically relevant economic headlines provided a suitable foundation for comparing framing across models and prompts.
Ten annotators participated in the study. Each was provided with detailed annotation guidelines (Appendix B) describing three ideological categories: Socialist, Neutral, and Capitalist. These categories were grounded in framing theory and focused on identifying underlying values, tonal cues, and thematic priorities, rather than lexical indicators. To prevent surface-level cueing, explicit ideological terminology (e.g., “socialism,” “capitalism,” “neutral”) was masked and replaced with “BLANK.” Annotators thus based their decisions on more nuanced features of framing such as tone, argumentation style, and thematic emphasis.
All examples used to define the ideological categories were drawn from authentic news sources representing distinct ideological leanings. This ensured annotator judgments were anchored in real-world framing conventions rather than potentially biased model outputs.
Annotators labeled a total of 500 articles generated by 7 large language models (LLMs), with each article having been prompted to reflect a specific ideological stance. Annotators were not informed of how the LLM was instructed to generate the articles and instead independently labeled the perceived ideological perspective of each article.
This setup enabled us to assess ideological alignment accuracy: whether each model successfully produced output consistent with its intended ideological prompt, as judged by independent human annotators.
Results
The annotation results enabled measurement of alignment accuracy for each LLM. For each article, we compared the annotators’ majority label to the original ideological prompt. Across all 500 articles and 7 models, the overall alignment accuracy was 75%, indicating that in three out of four cases, the model-generated output was perceived by annotators as reflecting the intended ideological perspective. Annotators noted that although some articles were clearly framed, others blended values or used ambiguous tones, and the nuances distinguishing ideological positions could often be subtle—underscoring the challenge of prompt adherence in generative models, particularly in ideologically sensitive contexts.
Table 5 shows that human annotation accuracy varied substantially across models (recall that we use the prompt/instruction given to the LLM as the “gold” label for computing annotation accuracy). The highest agreement was observed for Phi (87%) and Llama (84%), while Gemma and C4ai yielded the lowest agreement at 63% and 64%, respectively. These differences suggest that certain models may produce outputs with more discernible ideological cues or more consistent stylistic patterns, making them easier for annotators to classify. Conversely, models with lower agreement may produce more ambiguous or blended framings, complicating human interpretation.
Table 5: Human Annotation Accuracy per Model
Model Human Accuracy
C4ai 64%
Qwen 74%
Mistral 78%
GLM 80%
Gemma 63%
Phi 87%
Llama 84%
7.4 Model-Human Agreement on Neutral Article Detection
To evaluate how well LLMs detect bias, we conducted an experiment assessing agreement between model predictions and human judgments on identifying the neutral article within each set of three ideologically framed outputs. Unlike the alignment accuracy analysis, which measured whether a model’s output reflected its intended ideological prompt, this task required selecting the one neutral response from two biased alternatives.
The evaluation covered all available triplets in the dataset, comprising approximately 167 triplets (501 model-generated articles in total), with each triplet containing a Socialist, Capitalist, and Neutral version of the same headline. Ten annotators participated in the task. Where a triplet received multiple annotations, we derived the human label via majority vote to provide a stable reference point for model evaluation. [^3] See Table 9 in the Appendix for a breakdown of annotated triplets per model.
Models were prompted using the same annotation guidelines provided to annotators, ensuring consistent framing criteria. Both models and annotators independently selected the article judged to be most neutral in tone and framing.
To quantify agreement between model predictions and human judgments, we computed Cohen’s kappa scores, which account for chance agreement and provide a more robust reliability estimate than accuracy alone. Kappa values ranged from 0.40 (Command-R, Phi, C4ai, Mistral) to 0.47 (Qwen), indicating moderate agreement. Model accuracy on the neutral article identification task ranged from 68% to 85%, with several models—including Llama, Qwen, and Gemma—achieving accuracies above 80% (see Table 6).
Taken together, these results suggest that LLMs exhibit a reasonably strong ability to detect bias, achieving high accuracy and moderate agreement with human judgments. The moderate Cohen’s kappa scores (around 0.5) indicate that while models do not perfectly replicate human judgments, their outputs align with human consensus to a meaningful extent. However, these findings also suggest that the ability to detect bias does not equate to a preference for neutrality. Comparing these results with those in the right column of table 4 reveals that although models can distinguish between biased and neutral content, they often select biased content aligned with specific ideological positions. This indicates that even when LLMs recognize biased framing, they may still exhibit systematic preferences toward certain perspectives.
Table 6: Model Accuracy and Human Agreement on Neutral Article Detection
Model Model Accuracy Cohen’s Kappa Score
Llama 81% 0.43
Command-R 68% 0.40
Gemma 85% 0.45
Phi 78% 0.40
GLM 78% 0.45
Ministral 80% 0.40
Qwen 82% 0.47